Q14: If the electnic charge on C, is 300 uC, then the stored energy 01 b) o.063 c) 0.04 d) 0.023 e) 0.0025 (us 1 b) 100 )50 d) 200 e) 250 The charge on Ca (in uC) is:
Answer #1

14) we know the formula

V=Q/C=300/2=150\muF

and energy stored is=1/2CV2=1/2*2*10-6*(150*10-6)2=0.023J

so correct option is d)

15 ) we have the formula

V=Q/C=300/2=150\muF

so correct option is a)